# Chemical Signatures of AGB Mass Transfer in Gaia White Dwarf Companions

### v1

@@ 5.3 Ba enhancement in a NS companion?
We caution that the Gaia orbit of this source has not yet been validated (e.g. K. El-Badry et al., 2024d). It is possible that the implied secondary mass is overestimated, meaning that we cannot rule out the possibility of a massive WD secondary. We note that primary masses reported in the binary_masses table for metal-poor stars tend be overestimated (K. El-Badry et al., 2024b). Multi-epoch spectroscopic observations to measure and fit RVs will allow us to strengthen constraints on the orbit and nature of the companion.
@@
5.4 Carbon enhanced barium dwarfs
@@
We identify this absorption band as the G band of CH. This, along with the $\rm C_{2}$ Swan band (also visible in the spectra of the three stars; see top row of Figure 3), is among the only strong carbon features for metal-poor stars at these effective temperatures and therefore is commonly used in the study of CEMP stars (e.g. A. Goswami et al., 2006; M. Spite et al., 2013; C. J. Hansen et al., 2016; J. E. Norris & D. Yong, 2019). In the right column of Figure 9, we over-plot the spectrum of a known CEMP-s star, HE 0441-0652, from Ultraviolet and Visual Echelle Spectrograph (UVES; T. T. Hansen et al. 2016; L. Zhang et al. 2011; ESO proposal number 170.D-0010). The deep absorption of the three barium dwarfs match closely to those of the CEMP-s star. The fact that these features are absent in the star with no barium enhancement supports the idea that the both carbon and s-process elements originate from the same source – a WD companion that was formerly an AGB star. While we leave a more detailed analysis of the carbon abundance for future work, this suggests a smooth transition between the “barium dwarfs” and “CEMP-s” classifications (Section 6).
@@
6 Comparison to literature Ba dwarfs
@@
Figure 10: Barium dwarfs from this work and from the published literature, plotted in the space of: (A) eccentricity vs. orbital period, (B) primary vs. secondary mass, (C) [Ba/Fe] vs. [Fe/H], and (D) [Y/Fe] vs. [Fe/H] . We include the astrometric WD + barium dwarf binaries from P. Rekhi et al. (2025), barium stars/giants and dwarfs with spectroscopic orbits from (A. Jorissen et al., 2019) and A. Escorza et al. (2019) respectively, and CEMP-s stars (T. T. Hansen et al., 2016). Note that for systems with spectroscopic solutions, the secondary mass plotted is the minimum implied mass for an edge-on inclination. The systems from this work and P. Rekhi et al. (2025) reside in a similar region of these plots, but our additions extend the tails of the distributions. In particular, the most metal-poor systems in our sample bridge the gap between normal barium dwarfs and metal-poor CEMP-s stars.
@@
As discussed in P. Rekhi et al. (2025), the Gaia sample of WD + MS binaries contains barium dwarfs in a region of parameter space that was previously sparsely populated. In Figure 10, we compare several orbital and stellar properties of our systems containing barium dwarfs to those from P. Rekhi et al. (2025). Where available, we also include barium dwarfs and giants from the literature (A. Jorissen et al., 2019; A. Escorza et al., 2019) as well as CEMP-s stars (T. T. Hansen et al., 2016). Since the targets in our work and P. Rekhi et al. (2025) are all selected from the same underlying Gaia sample, on average, they occupy a similar region of the parameter space. However, several additional barium dwarfs from this work extend the tails of the distributions, strengthening previously observed trends and introducing interesting outliers.
@@
In particular, all systems from this work and P. Rekhi et al. (2025) have periods ranging from $\sim 100$ to 10001000\,d, which correspond to AU-scale separations, and most have eccentricities between 0 and 0.2. There is a positive relation between the component masses, as the astrometric triage technique requires more massive, and thus luminous, primaries to host more massive WDs (S. Shahaf et al., 2019, 2024). However, as highlighted in Section 5.3, we find two barium dwarfs around massive secondaries very close to or exceeding the Chandrasekhar limit, and these correspond to the two most eccentric systems with $e\sim 0.7$, making them consistent with having NS companions. In addition, there are two barium dwarfs around ultra-massive WD candidates with implied masses $\sim 1.2\,M_{\odot}$. It is worth mentioning that these systems have relatively low eccentricities ($\sim 0.2$), consistent with the bulk of the sample and unlike the two NS + MS binary candidates.
@@
There is an inverse trend between the barium abundance and metallicity. This trend is strengthened by the additions from our work, which extends the range of [Fe/H] from the P. Rekhi et al. (2025) sample down by more than 1 dex.
@@
Note that while the systems studied this work and P. Rekhi et al. (2024) were selected from the same underlying population from S. Shahaf et al. (2024), P. Rekhi et al. (2024) preferentially selected some targets based on having high eccentricities, and thus the different samples have heterogeneous selection functions. In addition, since the other literature barium dwarfs/stars were not selected based on their AMRFs (Section 2), they are on average more massive and luminous than those of our sample or P. Rekhi et al. (2025).
@@
As described in Section 5.4, we identified evidence of carbon enhancement in a few of the most metal-poor barium dwarfs in our sample. Indeed, these overlap in the space of [Ba/Fe] vs. [Fe/H] with the literature CEMP-s stars. There is no clear dividing line between the two populations, which suggests that they share a common physical origin.
@@
If both barium and carbon indeed originate from an AGB companion, it is natural to ask why carbon enhancement is not present in all barium dwarfs, but only the most metal-poor ones. One possible reason is the higher oxygen abundances in more metal-rich accretors. CEMP-s stars generally have high carbon to oxygen ratios ($\mathrm{[C/O]}\gtrsim 1$). In the presence of oxygen, carbon binds to it to form carbon monoxide (CO) which, unlike CH or C2, does not have strong spectral features in the optical (e.g. D. L. Lambert et al., 1986; T. C. Beers et al., 2007; C. R. Kennedy et al., 2011; Á. Skúladóttir et al., 2015; A. Gonneau et al., 2016). This may explain why CH and C2 absorption features are only found in the most metal-poor stars, where oxygen abundances are low. This, however, depends on CO molecules surviving in the atmospheres of these stars, which will not be true given high enough effective temperatures. Another possibility is that the accreted material is more carbon-rich at lower metallicities. Indeed, predictions of the carbon abundance, as well as the relative abundance of carbon to oxygen, in theoretical models of AGB stars increases with decreasing metallicity. This is shown in Section 7 (Figure 12).
7 What determines s-process enhancement?
@@
In Section 5.2, we studied how the s-process abundances of objects in our sample vary with binary parameters. Here, we discuss the physical origins of the trends, and lack thereof, by comparing to theoretical models of the donor AGB stars (i.e. WD progenitors) and the accreting primaries (i.e. MS stars).
In Figure 11, we plot predicted [Ba/Fe] for AGB star envelopes as a function of stellar radius for a range of initial stellar masses and three different metallicties ($Z$). These predictions are from the FRUITY database (S. Cristallo et al., 2009, 2015, 2016), which provides nucleosynthetic yields and abundances after each dredge-up episode on the AGB, obtained using a modified version of the evolutionary code FRANEC (A. Chieffi & O. Straniero, 1989; A. Chieffi et al., 1998; O. Straniero et al., 2006). A complementary database (ph-FRUITY) provides stellar parameters after every thermal pulse, which can be linked to the dredge-up episodes. We show models with $Z$ = 0.014 (solar), 0.002, and 0.00024, corresponding to [Fe/H] $\sim$ 0, -0.8, and -1.8 which roughly spans the metallicities of our observed systems.
[Ba/Fe] is predicted to increase as the AGB star evolves and its maximum radius grows. Thus, the expected enhancement in the accretor due to MT from an AGB donor depends on when MT occurs. Therefore, we may expect a correlation between measured [Ba/Fe] of the MS stars and the binaries’ initial separations. However, this evidently does not translate to a strong correlation with the present-day separations (Section 5.2; Figure 8). This is not unexpected, as MT can change the orbital separation in ways that may depend non-trivially on the binary parameters (see Section 7.1 for further discussion).
The predicted [Ba/Fe] also varies with $Z$. For stars of all masses, [Ba/Fe] generally increases with decreasing $Z$ at a fixed radius. This partially explains the trend between the observed s-process abundances of our barium dwarfs with their metallicities (Section 5.1; Figure 5). For masses between $\sim 2.0$ and $4.0\,M_{\odot}$, the trend is broken at early times (i.e. at smaller radii), which may add scatter to the observed trend.
@@
Additionally, in the left panel of Figure 12, we plot the predicted [C/H] ($\rm=[C/Fe]-[Fe/H]\approx[C/Fe]-log_{10}(Z)$) as a function of stellar radius and metallicity for a fixed initial mass of $2\,M_{\odot}$. [C/H] generally increases with decreasing metallicity, which we may expect to result in more carbon-rich accretors in metal-poor systems. In the right panel, we plot the predicted [C/O] which similarly shows an anti-correlation with metallicity, though more pronounced throughout the AGB evolution. If the carbon is indeed effectively locked up in carbon monoxide, this may be another important explanation for the CEMP stars (Section 6).
@@
7.1 Population modeling of Ba dwarf binaries
@@
We build on the results of N. Yamaguchi et al. (2025), who used a forward modeling approach to create a mock non-class I sample with properties that closely approximate the true S. Shahaf et al. (2024) sample. They start with a simulated population of Galactic zero-age binaries, then model the stellar evolution of each component as well as their orbital evolution due to binary interactions to obtain the present-day population of WD + MS binaries. In their fiducial model, the change in orbital separation due to MT from AGB donors for systems with initial accretor-to-donor mass ratios exceeding $\sim 0.4$ – which ultimately become the AU-scale WD + MS binaries that we study in this work – is calculated using the relations derived by G. E. Soberman et al. (1997) for fully non-conservative MT (i.e. $\beta=1$). Epoch astrometry is predicted for each present-day binary based on the Gaia scanning law which is then fit with astrometric models. A subset of these orbital solutions form the catalog of astrometric binaries which reproduces the key features of the true catalog published in Gaia DR3, suggesting that the selection function derived by this method is reasonable (K. El-Badry et al., 2024a). Finally, as done by S. Shahaf et al. (2024), non-class I systems are selected based on the AMRF. We refer readers to N. Yamaguchi et al. (2025) for an in-depth description of this process.
@@
In the top of row of Figure 14, we plot the predicted $\mathrm{Ba/Fe}]_{f}$ as a function of several orbital and stellar parameters for the simulated sample using the fiducial model of orbital evolution adopted by N. Yamaguchi et al. (2025). In total, fewer than 10% of the simulated MS stars have $\mathrm{Ba/Fe}]_{i}>0.25\,$dex. This fraction is primarily determined by the assumed initial separations of the binaries. As shown in Figure 11, [Ba/Fe] increases rapidly with the radius/age of the AGB donor. In the fiducial model, systems with initial Roche lobe radii $\lesssim 200\,R_{\odot}$ comprise the bulk of those that end up as AU-scale WD + MS binaries with $P_{\rm orb}<1000\,$d and thus enter our sample. AGB stars at these radii are in very early phases of s-process nucleosynthesis, and thus the accreted material is not yet strongly enriched in barium.
Conversely, if we adopt a different model such that systems with initially wider orbits experience more orbital shrinkage to form the AU-scale systems, MT would have begun when the AGB donors were more evolved, resulting in a larger fraction of barium dwarfs. As one example, if we continue to assume fully non-conservative MT but with some mass loss via wind from the donor ($A=5,\alpha=0.2,\beta=0.8$; G. E. Soberman et al. 1997), our systems would have experienced MT from AGB donors with radii as large as $\sim 400\,R_{\odot}$. This increases the predicted fraction of barium dwarfs to $\sim 30\,$%. We show the results of this model in the bottom row of Figure 14. Although the relation between initial to final separation is not one-to-one, there is a general trend of $\mathrm{Ba/Fe}]_{f}$ with the orbital period in the simulation, with a smaller fraction of barium dwarfs below $\sim 1\,$yr. We do not find evidence of such a trend in our observed systems (Figure 8). This likely reflects the fact that the true orbital evolution is more diverse and complex than assumed by our model, meaning there is more scatter in the relation between initial and final separations. Additionally, most classical barium stars/dwarfs are at periods >1000>1000\,d, so the fraction may still increase at longer periods.
@@
We do not attempt to further refine our simulation to reproduce the exact features of the observed systems and their measured s-process abundances. For this, we must consider the selection of our targets from the non-class I systems which is not homogeneous across the multiple observing programs. Moreover, there exists degeneracies in the orbital evolution predicted by the G. E. Soberman et al. (1997) relation for different choices of parameters which are difficult to disentangle. In addition, for simplicity, we only considered dilution in the context of the outermost convective envelope. However, in reality, non-convective mixing processes such as thermohaline mixing and atomic diffusion (e.g. J. Sevilla et al., 2022; A. Behmard et al., 2023) may have non-negligible effects, particularly as the convective layer becomes very thin at larger stellar masses and lower metallicities. It is also worth mentioning that the $\mathrm{[Ba/Fe]}_{i}=0$ assumption is a simplification as it has been shown to depend on [Fe/H] (e.g. P. E. Nissen & W. J. Schuster, 2011; H. Reggiani et al., 2017; A. M. Amarsi et al., 2020). However, we have confirmed that such a variation only reduces the total fraction of barium dwarfs in our simulated population by a few percent [^5] We tested a linear fit to data from H. Reggiani et al. (2017) and P. E. Nissen & W. J. Schuster (2011): $\mathrm{[Ba/Fe]}=-0.08625\mathrm{[Fe/H]}-0.14492$..
@@
This exercise suggests that a key factor in determining whether or not a star that has accreted material from an AGB donor becomes a barium dwarf is how evolved the donor was when interaction began. This is in turn determined by the initial separation of the binary. Dilution plays a minor role for systems in our sample, most of which contain MS stars with thin convective envelopes. As a result, the observed barium abundances do not place strong constraints on the MT efficiency. Overall, in the framework of this model, it is not surprising that many AU-scale WD + MS binaries are not enhanced in barium, simply because they interacted relatively early on the AGB. This may at least partially explain the fact that most previously discovered barium stars and related systems are in wider orbits than those of our sample (Figure 10). For the same reason, we may expect a larger fraction of barium dwarfs to be discovered in future Gaia data releases, which will be sensitive to binaries with longer orbital periods.
@@
Figure 14: Predicted barium abundance and barium dwarf fraction as a function of several orbital and stellar parameters for the simulated non-class I sample. For visual clarity, we plot a subset of 1500 systems. In magenta, we highlight the barium dwarfs, with [Ba/Fe] >0.25>0.25\,dex. In the top row are results using the fiducial model of orbital evolution via MT adopted by N. Yamaguchi et al. (2025), using relations from G. E. Soberman et al. (1997) with $\beta=1$ (i.e. fully non-conservative MT with mass loss via isotropic re-emission from the accretor). In the bottom row are results using the same relations, but with non-zero $\alpha$ (i.e. still fully non-conservative, but with some mass loss via winds from the donor). On average, the latter results in more orbital shrinkage, so that systems with initially wider separations and correspondingly more evolved AGB donors enter our sample. As s-process yields increase with time on the AGB, this leads to a larger fraction of barium dwarfs.
8 Conclusions

### v2

@@ V.3 Ba enhancement in a NS companion?
Additionally, this object shows evidence of carbon enhancement (Section V.4), which is also in tension with a massive SAGB origin. While dredge-up episodes in AGB stars are expected to bring up carbon to the outer layers and enrich them, this is thought to be avoided in SAGB stars, where the base of their convective envelopes reach sufficiently high temperatures to destroy carbon in a process known as “hot bottom burning” (e.g. Ventura et al., 2013; Doherty et al., 2014; Dell’Agli et al., 2017).
We caution that the Gaia orbit of this source has not yet been validated (e.g. El-Badry et al., 2024b). It is possible that the implied secondary mass is overestimated, meaning that we cannot rule out the possibility of a massive WD secondary. We note that primary masses reported in the binary_masses table for metal-poor stars tend be overestimated (El-Badry et al., 2024d). Multi-epoch spectroscopic observations to measure and fit RVs will allow us to strengthen constraints on the orbit and nature of the companion. If the abundances and companion mass are reliable, it is possible that the companion is a merger product of two lower-mass WDs (e.g. Hallakoun et al., 2026).
@@
V.4 Carbon enhanced barium dwarfs
@@
We identify this absorption band as the G band of CH. This, along with the $\rm C_{2}$ Swan band (also visible in the spectra of the three stars; see top row of Figure 3), is among the only strong carbon features for metal-poor stars at these effective temperatures and therefore is commonly used in the study of CEMP stars (e.g. Goswami et al., 2006; Spite et al., 2013; Hansen et al., 2016a; Norris and Yong, 2019). In the right column of Figure 9, we over-plot the spectrum of a known CEMP-s star, HE 0441-0652, from Ultraviolet and Visual Echelle Spectrograph (UVES; Hansen et al. 2016b; Zhang et al. 2011; ESO proposal number 170.D-0010). The deep absorption of the three barium dwarfs match closely to those of the CEMP-s star. The fact that these features are absent in the star with no barium enhancement supports the idea that the both carbon and s-process elements originate from the same source – a WD companion that was formerly an AGB star. While we leave a more detailed analysis of the carbon abundance for future work, this suggests a smooth transition between the “barium dwarfs” and “CEMP-s” classifications (Section VI).
@@
V.5 Sensitivity to the assumed stellar parameters and metallicity
In Section IV.2, we compared stellar parameters derived in this work to those from GALAH and found a systematic offset in [Fe/H] of $\sim 0.2\,$dex (Figure 2). While discrepancies between results from different methods are to be expected, here, we discuss their possible effects on the calculated s-process abundances.
If we naively take [Fe/H] values from GALAH as a reference for our measured [Ba/H], the calculated [Ba/Fe] (and similarly, [Y/Fe]) decreases. In this case, we find that out of the nine stars in our sample ultimately classified to be barium dwarfs that are also found in GALAH, six remain. All nine stars have FEROS spectra, with lower SNR than typical sources in our sample, so extrapolation of this fraction to the full sample is likely to yield a pessimistic estimate of the barium dwarf fraction.
On the other hand, if we fix all stellar parameters ($T_{\rm eff}$, [Fe/H], log(g)(g)) to those from GALAH and re-calculate our abundances, we find that the offsets in other parameters lead to a general increase in [Ba/H] that wins out over the higher metallicities. This means that there is a net increase in [Ba/Fe] and the number of stars classified to be barium dwarfs rises from 9 to 16. While both of these tests introduce internal inconsistencies in the method and therefore should not be taken at face value, they illustrate that the offsets here do not necessarily imply an overestimate in the fiducial barium dwarf fraction found in this work.
VI Comparison to literature Ba dwarfs
@@
Figure 10: Barium dwarfs from this work and from the published literature, plotted in the space of: (A) eccentricity vs. orbital period, (B) primary vs. secondary mass, (C) [Ba/Fe] vs. [Fe/H], and (D) [Y/Fe] vs. [Fe/H] . We include the astrometric WD + barium dwarf binaries from Rekhi et al. (2025), barium stars/giants and dwarfs with spectroscopic orbits from Jorissen et al. (2019) and Escorza et al. (2019) respectively, and CEMP-s stars (Hansen et al., 2016b). Note that for the barium stars and giants, the secondary mass plotted is the minimum implied mass for an edge-on inclination. Meanwhile, Hansen et al. (2016b) assumes a fixed primary mass of $0.8\,M_{\odot}$ and use this to derive the companion mass which we directly plot here. The systems from this work and Rekhi et al. (2025) reside in a similar region of these plots, but our additions extend the tails of the distributions. In particular, the most metal-poor systems in our sample bridge the gap between normal barium dwarfs and metal-poor CEMP-s stars.
@@
As discussed in Rekhi et al. (2025), the Gaia sample of WD + MS binaries contains barium dwarfs in a region of parameter space that was previously sparsely populated. In Figure 10, we compare several orbital and stellar properties of our systems containing barium dwarfs to those from Rekhi et al. (2025). Where available, we also include barium dwarfs and giants from the literature (Jorissen et al., 2019; Escorza et al., 2019) as well as CEMP-s stars (Hansen et al., 2016b). Since the targets in our work and Rekhi et al. (2025) are all selected from the same underlying Gaia sample, on average, they occupy a similar region of the parameter space. However, several additional barium dwarfs from this work extend the tails of the distributions, strengthening previously observed trends and introducing interesting outliers.
@@
In particular, all systems from this work and Rekhi et al. (2025) have periods ranging from $\sim 100$ to 10001000\,d, which correspond to AU-scale separations, and most have eccentricities between 0 and 0.2. There is a positive relation between the component masses, as the astrometric triage technique requires more massive, and thus luminous, primaries to host more massive WDs (Shahaf et al., 2019, 2024). However, as highlighted in Section V.3, we find two barium dwarfs around massive secondaries very close to or exceeding the Chandrasekhar limit, and these correspond to the two most eccentric systems with $e\sim 0.7$, making them consistent with having NS companions. In addition, there are two barium dwarfs around ultra-massive WD candidates with implied masses $\sim 1.2\,M_{\odot}$. It is worth mentioning that these systems have relatively low eccentricities ($\sim 0.2$), consistent with the bulk of the sample and unlike the two NS + MS binary candidates.
@@
There is an inverse trend between the barium abundance and metallicity. This trend is strengthened by the additions from our work, which extends the range of [Fe/H] from the Rekhi et al. (2025) sample down by more than 1 dex.
@@
Note that while the systems studied this work and Rekhi et al. (2024) were selected from the same underlying population from Shahaf et al. (2024), Rekhi et al. (2024) preferentially selected some targets based on having high eccentricities, and thus the different samples have heterogeneous selection functions. In addition, since the other literature barium dwarfs/stars were not selected based on their AMRFs (Section II), they are on average more massive and luminous than those of our sample or Rekhi et al. (2025).
@@
As described in Section V.4, we identified evidence of carbon enhancement in a few of the most metal-poor barium dwarfs in our sample. Indeed, these overlap in the space of [Ba/Fe] vs. [Fe/H] with the literature CEMP-s stars. There is no clear dividing line between the two populations, which suggests that they share a common physical origin.
@@
If both barium and carbon indeed originate from an AGB companion, it is natural to ask why carbon enhancement is not present in all barium dwarfs, but only the most metal-poor ones. While relative enhancement of any accreted element would be greater at lower metallicity, it is unclear why the strength of this effect varies between elements. One possible reason is the higher oxygen abundances in more metal-rich accretors. CEMP-s stars generally have high carbon to oxygen ratios ($\mathrm{[C/O]}\gtrsim 1$). In the presence of oxygen, carbon binds to it to form carbon monoxide (CO) which, unlike CH or C2, does not have strong spectral features in the optical (e.g. Lambert et al., 1986; Beers et al., 2007; Kennedy et al., 2011; Skúladóttir et al., 2015; Gonneau et al., 2016). This may explain why CH and C2 absorption features are only found in the most metal-poor stars, where oxygen abundances are low. This, however, depends on CO molecules surviving in the atmospheres of these stars, which will not be true given high enough effective temperatures. Another possibility is that the accreted material is more carbon-rich at lower metallicities. Indeed, predictions of the carbon abundance, as well as the relative abundance of carbon to oxygen, in theoretical models of AGB stars increases with decreasing metallicity. This is shown in Section VII (Figure 12).
VII What determines s-process enhancement?
@@
In Section V.2, we studied how the s-process abundances of objects in our sample vary with binary parameters. Here, we discuss the physical origins of the trends, and lack thereof, by comparing to theoretical models of the donor AGB stars (i.e. WD progenitors) and the accreting primaries (i.e. MS stars).
In Figure 11, we plot predicted [Ba/Fe] for AGB star envelopes as a function of stellar radius for a range of initial stellar masses and three different metallicties ($Z$). These predictions are from the FRUITY database (Cristallo et al., 2009, 2015, 2016), which provides nucleosynthetic yields and abundances after each dredge-up episode on the AGB, obtained using a modified version of the evolutionary code FRANEC (Chieffi and Straniero, 1989; Chieffi et al., 1998; Straniero et al., 2006). A complementary database (ph-FRUITY) provides stellar parameters after every thermal pulse, which can be linked to the dredge-up episodes. We show models with $Z$ = 0.014 (solar), 0.002, and 0.00024, corresponding to [Fe/H] $\sim$ 0, -0.8, and -1.8 which roughly spans the metallicities of our observed systems.
[Ba/Fe] is predicted to increase as the AGB star evolves and its maximum radius grows, corresponding to more dredge-up episodes and thermal pulses. Thus, the expected enhancement in the accretor due to MT from an AGB donor depends on when MT occurs. Therefore, we may expect a correlation between measured [Ba/Fe] of the MS stars and the binaries’ initial separations. However, this evidently does not translate to a strong correlation with the present-day separations (Section V.2; Figure 8). This is not unexpected, as MT can change the orbital separation in ways that may depend non-trivially on the binary parameters (see Section VII.1 for further discussion).
The predicted [Ba/Fe] also varies with $Z$. For stars of all masses, [Ba/Fe] generally increases with decreasing $Z$ at a fixed radius. This partially explains the trend between the observed s-process abundances of our barium dwarfs with their metallicities (Section V.1; Figure 5). For masses between $\sim 2.0$ and $4.0\,M_{\odot}$, the trend is broken at early times (i.e. at smaller radii), which may add scatter to the observed trend.
@@
Additionally, in the left panel of Figure 12, we plot the predicted [C/H] ($\rm=[C/Fe]-[Fe/H]\approx[C/Fe]-log_{10}(Z)$) as a function of stellar radius and metallicity for a fixed initial mass of $2\,M_{\odot}$. [C/H] generally increases with decreasing metallicity, which we may expect to result in more carbon-rich accretors in metal-poor systems. In the right panel, we plot the predicted [C/O] which similarly shows an anti-correlation with metallicity, though more pronounced throughout the AGB evolution. If the carbon is indeed effectively locked up in carbon monoxide, this may be another important explanation for the CEMP stars (Section VI).
@@
VII.1 Population modeling of Ba dwarf binaries
@@
We build on the results of Yamaguchi et al. (2025), who used a forward modeling approach to create a mock non-class I sample with properties that closely approximate the true Shahaf et al. (2024) sample. They start with a simulated population of Galactic zero-age binaries, then model the stellar evolution of each component as well as their orbital evolution due to binary interactions to obtain the present-day population of WD + MS binaries. In their fiducial model, the change in orbital separation due to MT from AGB donors for systems with initial accretor-to-donor mass ratios exceeding $\sim 0.4$ – which ultimately become the AU-scale WD + MS binaries that we study in this work – is calculated using the relations derived by Soberman et al. (1997) for fully non-conservative MT (i.e. $\beta=1$). Epoch astrometry is predicted for each present-day binary based on the Gaia scanning law which is then fit with astrometric models. A subset of these orbital solutions form the catalog of astrometric binaries which reproduces the key features of the true catalog published in Gaia DR3, suggesting that the selection function derived by this method is reasonable (El-Badry et al., 2024a). Finally, as done by Shahaf et al. (2024), non-class I systems are selected based on the AMRF. We refer readers to Yamaguchi et al. (2025) for an in-depth description of this process.
@@
In the top of row of Figure 14, we plot the predicted $\mathrm{Ba/Fe}]_{f}$ as a function of several orbital and stellar parameters for the simulated sample using the fiducial model of orbital evolution adopted by Yamaguchi et al. (2025). In total, fewer than 10% of the simulated MS stars have $\mathrm{Ba/Fe}]_{i}>0.25\,$dex. This fraction is primarily determined by the assumed initial separations of the binaries. As shown in Figure 11, [Ba/Fe] increases rapidly with the radius/age of the AGB donor. In the fiducial model, systems with initial Roche lobe radii $\lesssim 200\,R_{\odot}$ comprise the bulk of those that end up as AU-scale WD + MS binaries with $P_{\rm orb}<1000\,$d and thus enter our sample. AGB stars at these radii are in very early phases of s-process nucleosynthesis, and thus the accreted material is not yet strongly enriched in barium.
Conversely, if we adopt a different model such that systems with initially wider orbits experience more orbital shrinkage to form the AU-scale systems, MT would have begun when the AGB donors were more evolved, resulting in a larger fraction of barium dwarfs. As one example, if we continue to assume fully non-conservative MT but with some mass loss via wind from the donor ($A=5,\alpha=0.2,\beta=0.8$; Soberman et al. 1997), our systems would have experienced MT from AGB donors with radii as large as $\sim 400\,R_{\odot}$. This increases the predicted fraction of barium dwarfs to $\sim 30\,$%. We show the results of this model in the bottom row of Figure 14. Although the relation between initial to final separation is not one-to-one, there is a general trend of $\mathrm{Ba/Fe}]_{f}$ with the orbital period in the simulation, with a smaller fraction of barium dwarfs below $\sim 1\,$yr. We do not find evidence of such a trend in our observed systems (over-plotted with orange empty bars; also see Figure 8). This likely reflects the fact that the true orbital evolution is more diverse and complex than assumed by our model, meaning there is more scatter in the relation between initial and final separations. Additionally, most classical barium stars/dwarfs are at periods >1000>1000\,d, so the fraction may still increase at longer periods.
@@
We do not attempt to further refine our simulation to reproduce the exact features of the observed systems and their measured s-process abundances. For this, we must consider the selection of our targets from the non-class I systems which is not homogeneous across the multiple observing programs. Moreover, there exists degeneracies in the orbital evolution predicted by the Soberman et al. (1997) relation for different choices of parameters which are difficult to disentangle. In addition, for simplicity, we only considered dilution in the context of the outermost convective envelope. However, in reality, non-convective mixing processes such as thermohaline mixing and atomic diffusion (e.g. Sevilla et al., 2022; Behmard et al., 2023) may have non-negligible effects, particularly as the convective layer becomes very thin at larger stellar masses and lower metallicities. It is also worth mentioning that the $\mathrm{[Ba/Fe]}_{i}=0$ assumption is a simplification as it has been shown to depend on [Fe/H] (e.g. Nissen and Schuster, 2011; Reggiani et al., 2017; Amarsi et al., 2020). However, we have confirmed that such a variation only reduces the total fraction of barium dwarfs in our simulated population by a few percent [^5] We tested a linear fit to data from Reggiani et al. (2017) and Nissen and Schuster (2011): $\mathrm{[Ba/Fe]}=-0.08625\mathrm{[Fe/H]}-0.14492$..
@@
This exercise suggests that a key factor in determining whether or not a star that has accreted material from an AGB donor becomes a barium dwarf is how evolved the donor was when interaction began. This is in turn determined by the initial separation of the binary. Dilution plays a minor role for systems in our sample, most of which contain MS stars with thin convective envelopes. As a result, the observed barium abundances do not place strong constraints on the MT efficiency. Overall, in the framework of this model, it is not surprising that many AU-scale WD + MS binaries are not enhanced in barium, simply because they interacted relatively early on the AGB. This may at least partially explain the fact that most previously discovered barium stars and related systems are in wider orbits than those of our sample (Figure 10), though we caution that most previously known barium stars are giants, which have deeper convective envelopes than our targets and thus have experienced more dilution of accreted material. Future Gaia data releases will be sensitive to binaries with longer orbital periods, further testing how the barium star fraction varies with period and enabling a more direct comparison to literature barium star samples.
@@
Figure 14: Predicted barium abundance and barium dwarf fraction as a function of several orbital and stellar parameters for the simulated non-class I sample. For visual clarity, we plot a subset of 1500 systems. In magenta, we highlight the barium dwarfs, with [Ba/Fe] >0.25>0.25\,dex. In the top row are results using the fiducial model of orbital evolution via MT adopted by Yamaguchi et al. (2025), using relations from Soberman et al. (1997) with $\beta=1$ (i.e. fully non-conservative MT with mass loss via isotropic re-emission from the accretor). In the bottom row are results using the same relations, but with non-zero $\alpha$ (i.e. still fully non-conservative, but with some mass loss via winds from the donor). For reference, we over-plot the observed barium dwarf fractions in orange (Figure 8). On average, the latter model results in more orbital shrinkage, so that systems with initially wider separations and correspondingly more evolved AGB donors enter our sample. As s-process yields increase with time on the AGB, this leads to a larger fraction of barium dwarfs.
VIII Conclusions